Circular No. 5726

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET SHOEMAKER-LEVY (1993e)

Further precise positions have been reported as follows:

1993 UT α_{2000} δ_{2000} m_{1} Observer
Mar. 27.12479 12^{h}25^{m}17.^{s}77 −3^{o}55′32′′.3  McCrosky
27.14850 12 25 16.94 −3 55 24.6  _{”}
27.30417 12 25 12.37 −3 54 57.6 15 Tatum

R. E. McCrosky and C.-Y. Shao (Oak Ridge). 1.5-m reflector + CCD.

‘Meteoric’ appearance! ‘Train’ 41′′ long in p.a. 78^{o}-258^{o}. J. B. Tatum and D. D. Balam (Victoria). 0.25-m Schmidt.

The following parabolic orbital elements satisfy the nine observations and indicate minimum approach to Jupiter 0.31 AU on 1993 Mar. 30:

T = 1994 Aug. 3.441 TT  ω = 300.^{o}459
Ω = 344.563 }2000.0
q = 2.36996 AU  i = 2.322

The following representative elliptical orbital elements give the minimum approach to Jupiter as 0.04 AU on 1992 July 28:

T = 1988 Dec. 5.767 TT  ω = 43.^{o}253
e = 0.18714  Ω = 347.657 }2000.0
q = 3.91756 AU  i = 2.684

This ellipse puts the comet at r = 5.4 AU around Jupiter encounter with Jupiter-free values of q = 5.8 and 5.2 AU in 1979 and 1998, respectively. The ends of the Mar. 27 Oak Ridge train can be satisfied by varying the place in orbit by ±0.007 day at Jupiter encounter. Considering only its distance, the comet would have been as bright at its Feb. 1992 opposition as it is now, but splitting presumably occurred near Jupiter (as with P/Brooks 2 in 1886), and this is presumably to be associated with brightening.

1993 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Feb. 22 12^{h}39.^{m}75 − 5^{o}16.^{′}9 4.776 5.591 142.^{o}3 6.^{o}2 13.9
Mar. 4 12 36.04 − 4 57.1 4.694 5.598 153.3 4.6 13.8
14 12 31.65 − 4 32.5 4.640 5.604 164.4 2.7 13.8
24 12 26.84 − 4 04.6 4.616 5.611 175.6 0.8 13.8
Apr. 3 12 21.93 − 3 35.3 4.623 5.617 173.0 1.2 13.8
13 12 17.26 − 3 06.9 4.660 5.623 162.0 3.2 13.8
23 12 13.11 − 2 41.2 4.727 5.628 151.1 5.0 13.9

1993 March 27 Brian G. Marsden